A 25.00 mL sample of 0.379 M nitric acid (HNO3(aq)) is titrated with 12.00 mL of 0.573 M lithium hydroxide (LiOH(aq)). Determine the amount of [H+(aq)] remaining in the flask and pH of the solution after reaction
1
Expert's answer
2019-08-26T05:41:50-0400

"HNO_3(aq) + LiOH(aq)\rightarrow LiNO_3(aq) + H_2O(l)"

1) "n=c\times V"

"n(HNO_3) = 0.379 \frac{mol}{L} \times 25.00\times 10^{-3}L = 9.475\times10^{-3}mol"


"n(LiOH) = 0.573 \frac{mol}{L}\times12.00\times 10^{-3} L = 6.876\times 10^{-3}mol"


2) As mole ratio "n(HNO_3):n(LiOH) = 1:1" , then "n(HNO_3)" should be equal to "n(LiOH)"

We have "n(HNO_3)>n(LiOH)",

then LiOH is a limiting reactant.


3) "n(HNO_3)" consumed in the reaction is equal to "n(LiOH)"

"n(HNO_3)_{consumed} = 6.976\times10^{-3}mol"

"n(HNO_3)_{left} = 9.475\times10^{-3} - 6.876\times10^{-3} = 2.599\times 10 ^{-3}mol"


4) "HNO_3(aq)\rightarrow H^+(aq) + NO_3^-(aq)"

as 1 mole of "HNO_3" gives 1 mole of "H^+" then "n(H^+) = n(HNO_3)_{left} =2.599\times10^{-3} mol"

"c(H^+)=\frac{n(H^+)}{V} = \frac{2.599\times 10^{-3}mol}{25.00\times10^{-3}+12\times10^{-3}} = 7.024\times10^{-2} \frac{mol}{L}"


"pH =-log_{10}[H^+] -log_{10}7.024\times10^{-2} = 1.153"